I am trying to come up with an algebraic equation to determine the cost of two refrigerators with good energy ratings. I found one with that will cost .1143 cents per kWh and it boasts that it uses 486 kWh per year. I want to compare it with another one I found that will cost .1143 cents per kWh. This one uses 526 kWh per year. If the purchase price of the first refrigerator is $950 and the second is $875, which one will be the better buy after 10 years?
I'm sorry for the long question, I hope someone out there can give me a simple algebraic equation for putting this all together. I'd like to use it again for other appliances in the future.

Answer by mananth(16949)

First refrigerator cost + power cost for ten years
General Equation:
Let the number of years run be x
(Cx1)=950+0.1143*486*x
(Cx)=950 + 55.55x
Second refrigerator
(Cx2)=875+0.1143*526*x
(Cx2)=875+60.12x
...
Plug the value of x=10 for both equations to find which is a better deal.
(Cx1)= $1505.5
(Cx2)= $ 1476.2
..
You decide
..
m.ananth@hotmail.ca